# Safe Explicable Planning

### v3

@@ Introduction
In our Safe Explicable Planning (SEP) approach, we build upon the following assumptions to focus on the planning challenges. First, we assume that the agent has access to its model ($\mathcal{M}_{R}$) and the human’s belief of its model ($\mathcal{M}_{R}^{H}$), or simply, the human’s model. A similar assumption has been made in prior research on explicable planning  (Kulkarni et al. 2016; Hanni and Zhang 2021) and explainable decision-making  (Chakraborti et al. 2019). In practice, the human’s model may be provided by experts or acquired from human feedback, which has been explored in previous studies (Christiano et al. 2017; Ibarz et al. 2018; Holmes et al. 2004; Juba and Stern 2022). Second, we assume that the human is a rational observer, i.e., a behavior with a higher expected return in the human’s model is more expected. Hence, the most expected behavior can be generated by computing the optimal behavior in the human’s model. This assumption allows us to simplify the problem; from maximizing explicability to maximizing the expected return of a policy under the human’s model modeled as an MDP. Such an assumption of human rationality is a common simplification in cognitive science and artificial intelligence research, such as in  (Baker, Saxe, and Tenenbaum 2011).
We formulate Safe Explicable Planning (SEP) under MDP by defining the objective as maximizing the expected return in the human’s model, subject to a constraint in the agent’s model. This problem formulation generalizes the consideration of multiple objectives  (Marler and Arora 2004) to also consider multiple domain models. The solution to this problem yields a Pareto set of policies for which exact solvers are often intractable. To address this challenge, we propose an action-pruning technique to reduce the policy space significantly. Subsequently, we introduce a novel tree search method that efficiently explores the remaining policies to identify the Pareto set. We formally prove that this search method is sound and complete. Additionally, we introduce a greedy search method for situations where any policy from the Pareto set suffices. Finally, we devise approximate solutions for both search methods using state aggregation, addressing scalability in complex domains. We evaluate our methods across several domains via simulation and physical robot experiments, demonstrating their effectiveness for SEP. Furthermore, we conduct ablation studies to analyze the benefits of our pruning techniques, validating their effectiveness in reducing computational costs while generating the desired behaviors.
@@
Our problem formulation of SEP shares some key features with the constrained-criterion-based formulation of safe reinforcement learning (RL) (Garcıa and Fernández 2015), which is inherently a Constrained MDP (CMDP) (Altman 2021). Similar problem formulations have been proposed for continuous spaces and applied to risk-bounded motion planning (Huang et al. 2019). In these prior works, safety is encoded by constraining the expected cost under some designated cost function while maximizing the agent’s reward function under the same model. In SEP, similarly, safety is encoded by constraining the expected return under the agent’s reward function. SEP operates under the assumption that safety directly correlates with the expected return in the agent’s model, following the intuition that unsafe behaviors would result in low returns. Our formulation can readily accommodate a CMDP (with a single constraint) by aligning the two different models (except for the reward functions) and substituting the robot’s reward function in the safety constraint with the cost function.
@@
The objective considered in SEP also bears a similarity to that in Multi-Objective Markov Decision Processes (MOMDP) (White 1982), as SEP must consider the expected return under both the agent’s and human’s model. MOMDPs, introduced for multiple objectives under the same MDP (refer to the review paper by (Roijers et al. 2013)), typically aim to optimize a vector of expected returns for those objectives to derive a Pareto set of policies or to derive a single policy through linear scalarization of those objectives. Approaches, including but not limited to  (Wakuta and Togawa 1998; Russell and Zimdars 2003; Barrett and Narayanan 2008; Van Moffaert, Drugan, and Nowé 2013), are examples of these methods. However, if multiple domain dynamics models are present, MOMDP methods must produce multiple vectors of expected returns, one for each different model. Optimizing these vectors simultaneously poses a significantly greater challenge than optimizing a single vector in traditional MOMDPs.
@@
We work with the set of all stationary deterministic policies $\Pi$, given by $\forall\pi\in\Pi$, $\pi:\mathcal{S}\mapsto\mathcal{A}$. An optimal agent’s policy maximizes the expected return in the agent’s model and is given by $\pi^{*}=\operatorname{arg\,max}_{\pi}\mathbb{E}_{\mathcal{T}_{R}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma_{R}^{t}r_{R}(t)\right]$. We define a safe behavior as any behavior with a return within a bound of the optimal agent’s return. Similar criteria have been used in safe RL (Garcıa and Fernández 2015; Moldovan and Abbeel 2012). More formally, a policy $\pi$ is considered safe if its return satisfies the following condition:
@@
In prior work on explicable planning, the objective is to maximize a weighted sum of the return in the agent’s model and an explicability metric. Explicability metric has been defined, for example, via plan distances (Kulkarni et al. 2016) in deterministic domains and KL divergence between trajectory distributions (Gong and Zhang 2022) in stochastic domains. In our work, we define the explicability metric simply as the return from the human’s model $\mathcal{M}_{R}^{H}$. Given that the human user generates expectations from $\mathcal{M}_{R}^{H}$, this assumes a rational human observer: the higher the return in the human’s model, the more expected the policy is.
@@
Safe Explicable Planning (SEP), given by $\mathcal{P}_{\mathcal{E}}=\langle\mathcal{M}_{R},\mathcal{M}_{R}^{H},\delta\rangle$, is the problem to search for a policy that maximizes the return in $\mathcal{M}_{R}^{H}$ subject to a constraint on the return in $\mathcal{M}_{R}$ under any state, or more formally:
@@
Requiring the constraint and optimality above to hold under any state introduces a Pareto set of optimal policies where no policies in this set are strictly dominated by any policy. Briefly, a policy $\pi_{1}$ strictly dominates another policy $\pi_{2}$ if its state values are no smaller in any state, and larger in at least one state. More formally, denote such a relationship as $\pi_{1}\succ\pi_{2}$, which holds if $\forall s\in\mathcal{S}$ $[V_{\mathcal{M}_{R}^{H}}^{\pi_{1}}(s)\geq V_{\mathcal{M}_{R}^{H}}^{\pi_{2}}(s)]$wedge $\exists s^{\prime}\in\mathcal{S}$ $[V_{\mathcal{M}_{R}^{H}}^{\pi_{1}}(s^{\prime})>V_{\mathcal{M}_{R}^{H}}^{\pi_{2}}(s^{\prime})]$. The Pareto set $\Pi^{*}_{\mathcal{E}}$ is then given by:
@@
In this section, we motivate and discuss our solution methods for SEP. Given the large policy space to search for, we first discuss a technique to reduce the policy space. Since any policy $\Pi_{\delta}$ may be in the Pareto set, it necessitates the expansion of all policies in $\Pi_{\delta}$ We propose an exact method that selectively expands policies in $\Pi_{\delta}$ to determine the Pareto set $\Pi^{*}_{\mathcal{E}}$ Additionally, we discuss a greedy method that expands only a subset of policies in $\Pi_{\delta}$ , returning a single policy in $\Pi^{*}_{\mathcal{E}}$ Finally, we propose approximate solutions via state aggregation, using handcrafted features, to condition similar states to choose the same actions to further scalability in complex domains. Complete proofs, evaluation details, and additional discussions are presented in the full version of this paper (Hanni, Boateng, and Zhang 2023).
@@
To determine $\Pi^{*}_{\mathcal{E}}$, intuitively, we can evaluate every policy in $\widetilde{\Pi}$. However, this would be impractical and proves to be unnecessary. A more efficient strategy involves further reducing $\widetilde{\Pi}$ by expanding policies in a specific order. There are two possible search strategies to explore. First, consider initializing the search to the optimal policy in the human’s model and perform policy improvement under the agent’s objective until the bound is satisfied. Alternatively, consider initializing search to the optimal policy in the agent’s model and perform policy decrement under the agent’s objective while simultaneously identifying improving policies under the human’s objective, until the bound is violated. While the first search strategy is simpler it can lead to missed policies in $\Pi^{*}_{\mathcal{E}}$, hence we choose the latter option in our work.
@@
$\sum_{s^{\prime}}\mathcal{T}_{R}(s,a,s^{\prime})V_{\mathcal{M}_{R}}^{\pi}(s^{\prime})\leq\sum_{s^{\prime}}\mathcal{T}_{R}(s,\pi(s),s^{\prime})V_{\mathcal{M}_{R}}^{\pi}(s^{\prime}).$ (5)
@@
$\sum_{s^{\prime}}\mathcal{T}_{R}^{H}(s,a,s^{\prime})V_{\mathcal{M}_{R}^{H}}^{\pi}(s^{\prime})\geq\sum_{s^{\prime}}\mathcal{T}_{R}^{H}(s,\pi(s),s^{\prime})V_{\mathcal{M}_{R}^{H}}^{\pi}(s^{\prime}),$ (6)
@@
Bound Selection: In our approach, we assume the bound is specified by the designer, based on experience. However, it can often be estimated based on the domain. For instance, consider one of the cliff worlds depicted in Fig. 5 (see below). The optimal return in the agent’s model is $94$ (i.e., moving along the edge of the cliff to the goal), while the return of the trajectory with the longest detour (i.e., staying as far away from the edge as possible) without falling off the cliff is $90$, discount notwithstanding. As unsafe behaviors yield significantly lower returns (in $\mathcal{M}_{R}$) than the detour the safety bound can be set to $90/94{=}0.957$, subject to adjustment. Further analyses will be deferred to future work.
@@ Results:
3)3)3 ) Behavior Comparison in Wumpus World: The results are shown in Fig. 3. Following the optimal policy in the agent’s model ($\mathcal{M}_{R}$), the agent collected both coins while staying near the wumpus before exiting, as shown in Fig. 3(a). Following the optimal policy in the human’s modelexpected policy (from $\mathcal{M}_{R}^{H}$), the agent avoided getting close to the wumpus and collected a single coin before exiting, as shown in Fig. 3(b). When applying SEP under the bound $\delta=0.90$, PDT+ returns a large Pareto set (see Tab. 1). To select from $\Pi^{*}_{\mathcal{E}}$, we score policies based on the average distance between the agent and the wumpus throughout the most likely trajectory. The trajectory from the policy with the highest score is shown in Fig. 3(c): we can observe that the agent managed to collect both coins while maintaining a cautious distance from the wumpus, albeit taking a longer path, which is more explicable than the agent’s optimal behavior in Fig. 3(a) and simultaneously more efficient than the human’s expectation in Fig. 3(b). Fig. 3(d) showcases the behavior obtained by PAG+, which also maintains a cautious distance from the wumpus, for the most part.

### v4

@@ Introduction
In our Safe Explicable Planning (SEP) approach, we build upon the following assumptions to focus on the planning challenges. First, we assume that the agent has access to its model ($\mathcal{M}_{R}$) and the human’s belief of its model ($\mathcal{M}_{R}^{H}$), or simply, the human’s model. A similar assumption has been made in prior research on explicable planning  (Kulkarni et al. 2016; Hanni and Zhang 2021) and explainable decision-making  (Chakraborti et al. 2019). In practice, the human’s model may be provided by experts or acquired from human feedback, which has been explored in previous studies (Christiano et al. 2017; Ibarz et al. 2018; Holmes et al. 2004; Juba and Stern 2022). Second, we assume that the human is a rational observer, i.e., a behavior with a higher expected return in the human’s model is more expected. Hence, the most expected behavior can be generated by computing the optimal behavior in the human’s model. This assumption allows us to equate the problem of maximizing explicability to maximizing the expected return of a policy under the human’s model (that is modeled as an MDP). Such an assumption of human rationality is a common simplification in cognitive science and artificial intelligence research, such as in  (Baker, Saxe, and Tenenbaum 2011).
We formulate Safe Explicable Planning (SEP) under MDP by defining the objective as maximizing the expected return in the human’s model, subject to a constraint in the agent’s model. This problem formulation generalizes the consideration of multiple objectives (White 1982) to also consider multiple domain models. The solution to this problem yields a Pareto set of policies for which exact solvers are often intractable. To address this challenge, we propose an action-pruning technique to reduce the policy space significantly. Subsequently, we introduce a novel tree search method that efficiently explores the remaining policies to identify the Pareto set. We formally prove that this search method is sound and complete. Additionally, we introduce a greedy search method for situations where any policy from the Pareto set suffices. Finally, we devise approximate solutions for both search methods using state aggregation, addressing scalability in large domains. We evaluate our methods across several domains via simulation and physical robot experiments, demonstrating their effectiveness for SEP. Furthermore, we conduct ablation studies to analyze the benefits of our pruning techniques, validating their effectiveness in reducing computational costs while generating the desired behaviors.
@@
Our problem formulation of SEP shares some key features with the constrained-criterion-based formulation of safe reinforcement learning (RL) (Garcıa and Fernández 2015), which is inherently a Constrained Markov Decision Process (CMDP) (Altman 2021). Similar problem formulations have been proposed for continuous spaces and applied to risk-bounded motion planning (Huang et al. 2019). In these prior works, safety is encoded by constraining the expected cost under some designated cost function while maximizing the agent’s reward function under the same model. In SEP, similarly, safety is encoded by constraining the expected return under the agent’s reward function. SEP operates under the assumption that safety directly correlates with the expected return in the agent’s model, following the intuition that unsafe behaviors would result in low returns. Our formulation can readily accommodate a CMDP (with a single constraint) by aligning the two different models (except for the reward functions) and substituting the robot’s reward function in the safety constraint with the cost function.
@@
The objective considered in SEP also bears a similarity to that in Multi-Objective Markov Decision Processes (MOMDP) (White 1982), as SEP must consider the expected return under both the agent’s and human’s model. MOMDPs, introduced for multiple objectives under the same MDP (refer to the review paper by (Roijers et al. 2013)), typically aim to optimize a vector of expected returns for those objectives to derive a Pareto set of policies or to derive a single policy through linear scalarization of those objectives. Approaches, including but not limited to  (Wakuta and Togawa 1998; Russell and Zimdars 2003; Barrett and Narayanan 2008; Van Moffaert, Drugan, and Nowé 2013), are examples of these methods. However, to handle different models, MOMDP methods must produce multiple vectors of expected returns, each derived for a different model due to the difference in domain dynamics. Optimizing these vectors simultaneously poses a significantly greater challenge than optimizing a single vector in traditional MOMDPs.
@@
We work with the set of all stationary deterministic policies $\Pi$, where $\forall\pi\in\Pi$, $\pi:\mathcal{S}\mapsto\mathcal{A}$. An agent’s optimal policy maximizes the expected return in the agent’s model and is given by $\pi^{*}=\operatorname{arg\,max}_{\pi}\mathbb{E}_{\mathcal{T}_{R}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma_{R}^{t}r_{R}(t)\right]$. We define a safe behavior as any behavior with a return within a bound of the agent’s optimal return. Similar criteria have been used in safe RL (Garcıa and Fernández 2015; Moldovan and Abbeel 2012). More formally, a policy $\pi$ is considered safe or feasible if its return satisfies the following condition:
@@
In prior work on explicable planning, the objective is to maximize a weighted sum of the return in the agent’s model and an explicability metric. Explicability metric has been defined, for example, via plan distances (Kulkarni et al. 2016) in deterministic domains and KL divergence between trajectory distributions (Gong and Zhang 2022) in stochastic domains. In our work, we define the explicability metric simply as the return in the human’s model $\mathcal{M}_{R}^{H}$. Given that the human user generates expectations from $\mathcal{M}_{R}^{H}$, this assumes a rational human observer: the higher the return in the human’s model, the more expected the policy is.
@@
Safe Explicable Planning (SEP), given by $\mathcal{P}_{\mathcal{E}}=\langle\mathcal{M}_{R},\mathcal{M}_{R}^{H},\delta\rangle$, is the problem to search for a policy that maximizes the return in $\mathcal{M}_{R}^{H}$ subject to a constraint on the return in $\mathcal{M}_{R}$ under any state, or formally:
@@
The maximization of the expected return above across all states introduces a Pareto set of optimal policies where no policies in this set are strictly dominated by any feasible policy. Briefly, a policy $\pi_{1}$ strictly dominates another policy $\pi_{2}$ if its state values are no smaller in any state, and larger in at least one state. Formally, we denote such a relationship as $\pi_{1}\succ\pi_{2}$, which holds if $\forall s\in\mathcal{S}$ $[V_{\mathcal{M}_{R}^{H}}^{\pi_{1}}(s)\geq V_{\mathcal{M}_{R}^{H}}^{\pi_{2}}(s)]$wedge $\exists s^{\prime}\in\mathcal{S}$ $[V_{\mathcal{M}_{R}^{H}}^{\pi_{1}}(s^{\prime})>V_{\mathcal{M}_{R}^{H}}^{\pi_{2}}(s^{\prime})]$. The Pareto set $\Pi^{*}_{\mathcal{E}}$ is then given by:
@@
In this section, we motivate and discuss our solution methods for SEP. Given the large policy space to search for, we first discuss a technique to reduce the policy space. Since any policy $\Pi_{\delta}$ may be in the Pareto set, it necessitates the expansion of all policies in $\Pi_{\delta}$ We propose an exact method that selectively expands policies in $\Pi_{\delta}$ to determine the Pareto set $\Pi^{*}_{\mathcal{E}}$ Additionally, we discuss a greedy method that expands only a subset of policies in $\Pi_{\delta}$ , returning a single policy in $\Pi^{*}_{\mathcal{E}}$ Finally, we propose approximate solutions via state aggregation, using handcrafted features, to condition similar states to choose the same actions to further scalability in large domains. Complete proofs, evaluation details, and additional discussions are presented in the full version of this paper (Hanni, Boateng, and Zhang 2024).
@@
To determine $\Pi^{*}_{\mathcal{E}}$, intuitively, we can evaluate every policy in $\widetilde{\Pi}$. However, this would be impractical and proves to be unnecessary. A more efficient strategy involves further reducing $\widetilde{\Pi}$ by expanding policies in a specific order. There are two possible search strategies to explore. First, consider initializing the search to the optimal policy in the human’s model and perform policy improvement under the agent’s objective until the bound is satisfied. Alternatively, consider initializing the search to the optimal policy in the agent’s model and perform policy decrement under the agent’s objective while simultaneously identifying better policies under the human’s objective, until the bound is violated. While the first search strategy is simpler it can lead to missed policies in $\Pi^{*}_{\mathcal{E}}$, hence we choose the latter option in our work.
@@
$Q_{\mathcal{M}_{R}}^{\pi}(s,a)\leq Q_{\mathcal{M}_{R}}^{\pi}(s,\pi(s)).$ (5)
@@
$Q_{\mathcal{M}_{R}^{H}}^{\pi}(s,a)\geq Q_{\mathcal{M}_{R}^{H}}^{\pi}(s,\pi(s)),$ (6)
@@
Bound Selection: In our approach, we assume the bound is specified by the designer, based on experience. However, it can often be estimated based on the domain. For instance, consider one of the cliff worlds depicted in Fig. 5 (see below). The optimal return in the agent’s model is $94$ (i.e., moving along the edge of the cliff to the goal), while the return of the trajectory with the longest detour (i.e., staying as far away from the edge as possible) without falling off the cliff is $90$, discount notwithstanding. As unsafe behaviors yield significantly lower returns (in $\mathcal{M}_{R}$) than the detour, the safety bound can be set to $90/94{=}0.957$, subject to adjustment. Further analyses will be deferred to future work.
@@ Results:
3)3)3 ) Behavior Comparison in Wumpus World: The results are shown in Fig. 3. Following the optimal policy in the agent’s model ($\mathcal{M}_{R}$), the agent collected both coins while staying near the wumpus before exiting, as shown in Fig. 3(a). Following the optimal policy in the human’s model ($\mathcal{M}_{R}^{H}$), the agent avoided getting close to the wumpus and collected a single coin before exiting, as shown in Fig. 3(b). When applying SEP under the bound $\delta=0.90$, PDT+ returns a large Pareto set (see Tab. 1). To select from $\Pi^{*}_{\mathcal{E}}$, we score policies based on the average distance between the agent and the wumpus throughout the most likely trajectory. The trajectory from the policy with the highest score is shown in Fig. 3(c): we can observe that the agent managed to collect both coins while maintaining a cautious distance from the wumpus, albeit taking a longer path, which is more explicable than the agent’s optimal behavior in Fig. 3(a) and simultaneously more efficient than the human’s expectation in Fig. 3(b). Fig. 3(d) showcases the behavior obtained by PAG+, which also maintains a cautious distance from the wumpus, for the most part.